OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, find the limit. equation in comments

OpenStudy (anonymous):

\[a_n = \frac{ n^2 }{ \sqrt{n^3+3n} }\]

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} a_n = ?\]

OpenStudy (anonymous):

DNE